# Dependability of UAV-Based Networks and Computing Systems: A Survey

### v1

@@ Acknowledgements.
- Wang et al. (2019d) Peng Wang, Jiaxin Zhang, Xing Zhang, Zhi Yan, Barry G Evans, and Wenbo Wang. 2019d. Convergence of satellite and terrestrial networks: A comprehensive survey. IEEE access 8 (2019), 5550–5588.
- Wang et al. (2024d) Qixu Wang, Peng Xiao, Xiang Li, Yunxiang Qiu, Tao Zheng, and Zhiguang Qin. 2024d. Taas: Trust assessment as a service for secure communication of green edge-assisted UAV network. Peer Peer Netw. Appl. 17, 4 (July 2024), 2221–2236.
@@
- Wang et al. (2021g) Rui Wang, Xiangping Bryce Zhai, Yunlong Zhao, and Xuedong Zhao. 2021g. Delivery optimization for unmanned aerial vehicles based on minimum cost maximum flow with limited battery capacity. In Lecture Notes in Computer Science. Springer International Publishing, Cham, 77–85.
@@
- Wang et al. (2019b) Wenbo Wang, Niloofar Okati, Islam Tanash, Taneli Riihonen, and Elena-Simona Lohan. 2019b. Location-based beamforming architecture for efficient farming applications with drones. In 2019 International Conference on Localization and GNSS (ICL-GNSS) (Nuremberg, Germany). IEEE.
@@
- Wang et al. (2021c) Xiaoding Wang, Hui Lin, Hongyan Zhang, Dejun Miao, Qinyang Miao, and Wenxin Liu. 2021c. Intelligent drone-assisted fault diagnosis for B5G-enabled space-air-ground-space networks. IEEE Trans. Netw. Sci. Eng. 8, 4 (Oct. 2021), 2849–2860.
- Wang et al. (2021d) Xiaoding Wang, Hui Lin, Hongyan Zhang, Dejun Miao, Qinyang Miao, and Wenxin Liu. 2021d. Intelligent drone-assisted fault diagnosis for B5G-enabled space-air-ground-space networks. IEEE Trans. Netw. Sci. Eng. 8, 4 (Oct. 2021), 2849–2860.
@@
- Wang et al. (2021b) Yuhang Wang, Ying He, and Minhui Dong. 2021b. Resource allocation in vehicular networks with multi-UAV served edge computing. In 2021 IEEE 29th International Conference on Network Protocols (ICNP) (Dallas, TX, USA). IEEE.
@@
- Wang et al. (2022d) Yuntao Wang, Zhou Su, Tom H Luan, Ruidong Li, and Kuan Zhang. 2022d. Federated learning with fair incentives and robust aggregation for UAV-aided crowdsensing. IEEE Trans. Netw. Sci. Eng. 9, 5 (Sept. 2022), 3179–3196.
- Wang et al. (2021e) Yuntao Wang, Zhou Su, Ning Zhang, and Dongfeng Fang. 2021e. Disaster relief wireless networks: Challenges and solutions. IEEE Wirel. Commun. 28, 5 (Oct. 2021), 148–155.
@@
- Wang et al. (2020d) Yangang Wang, Hai Wang, and Xianglin Wei. 2020d. Energy-efficient UAV deployment and task scheduling in multi-UAV edge computing. In 2020 International Conference on Wireless Communications and Signal Processing (WCSP) (Nanjing). IEEE.
@@
- Wang et al. (2024c) Yangang Wang, Xianglin Wei, Hai Wang, Yongyang Hu, Kuang Zhao, and Jianhua Fan. 2024c. Distributed matching theory-based task re-allocating for heterogeneous multi-UAV edge computing. China Commun. 21, 1 (Jan. 2024), 260–278.
- Wang et al. (2024e) Yiqian Wang, Jie Zhu, Haiping Huang, and Fu Xiao. 2024e. Bi-objective ant colony optimization for trajectory planning and task offloading in UAV-assisted MEC systems. IEEE Trans. Mob. Comput. 23, 12 (Dec. 2024), 12360–12377.
@@
- Wang et al. (2019c) Zhifang Wang, Jianguo Yu, Shangjing Lin, Junguo Dong, and Zhen Yu. 2019c. Distributed robust adaptive fault-tolerant mechanism for quadrotor UAV real-time wireless network systems with random delay and packet loss. IEEE Access 7 (2019), 134055–134062.
- Wang et al. (2021f) Zhifang Wang, Jianguo Yu, Shangjing Lin, Junguo Dong, and Zheng Yu. 2021f. Distributed robust H adaptive fault-tolerant control of amorphous and flat air-to-ground wireless self-organizing network system. Assem. Autom. 41, 6 (Nov. 2021), 641–658.
@@
- Zeng et al. (2024c) Kai Zeng, Xin Li, and Tao Shen. 2024c. Energy-stabilized computing offloading algorithm for UAVs with energy harvesting. IEEE Internet Things J. 11, 4 (Feb. 2024), 6020–6031.
@@
- Zhang et al. (2018a) Long Zhang, Zhen Zhao, Qiwu Wu, Hui Zhao, Haitao Xu, and Xiaobo Wu. 2018a. Energy-aware dynamic resource allocation in UAV assisted mobile edge computing over social internet of vehicles. IEEE Access 6 (2018), 56700–56715.
- Zhang et al. (2018b) Long Zhang, Zhen Zhao, Qiwu Wu, Hui Zhao, Haitao Xu, and Xiaobo Wu. 2018b. Energy-aware dynamic resource allocation in UAV assisted mobile edge computing over social internet of vehicles. IEEE Access 6 (2018), 56700–56715.
- Zhang et al. (2023e) Peng Zhang, Hui Tian, Hao Luo, Xingwang Li, and Gaofeng Nie. 2023e. A hybrid fast inference approach with distributed neural networks for edge computing enabled UAV swarm. Phys. Commun. 60, 102129 (Oct. 2023), 102129.
@@
- Zhang et al. (2023d) Xuhui Zhang, Wenchao Liu, Huijun Xing, Zhenzhen Jin, Weilin Zang, Shuqiang Wang, Yanyan Shen, and Liang Xue. 2023d. Learning to hybrid offload in space-air-ground integrated mobile edge computing for IoT networks. In 2023 IEEE 13th International Conference on CYBER Technology in Automation, Control, and Intelligent Systems (CYBER) (Qinhuangdao, China). IEEE, 836–841.
@@
- Zhang et al. (2023c) Yafeng Zhang, Yongzheng Gan, Chunlin Li, Chunping Deng, and Youlong Luo. 2023c. Primary node selection based on node reputation evaluation for PBFT in UAV-assisted MEC environment. Wirel. Netw. 29, 8 (Nov. 2023), 3515–3539.
@@
- Zhang et al. (2023f) Ying Zhang, Supu Xiu, Yiqing Cai, and Pengshan Ren. 2023f. Scheduling of graph neural network and Markov based UAV mobile edge computing networks. Phys. Commun. 60, 102160 (Oct. 2023), 102160.
@@
- Zheng et al. (2024a) Chunxiang Zheng, Kai Pan, Jiadong Dong, Lin Chen, Qinghu Guo, Shunfeng Wu, Hanyun Luo, and Xiaolin Zhang. 2024a. Multi-agent collaborative optimization of UAV trajectory and latency-aware DAG task offloading in UAV-assisted MEC. IEEE Access 12 (2024), 42521–42534.
- Zheng et al. (2024b) Chunxiang Zheng, Kai Pan, Jiadong Dong, Lin Chen, Qinghu Guo, Shunfeng Wu, Hanyun Luo, and Xiaolin Zhang. 2024b. Multi-agent collaborative optimization of UAV trajectory and latency-aware DAG task offloading in UAV-assisted MEC. IEEE Access 12 (2024), 42521–42534.
@@
- Zhou et al. (2024d) Xunkuai Zhou, Guidong Yang, Yizhou Chen, Li Li, and Ben M Chen. 2024d. VDTNet: A high-performance visual network for detecting and tracking of intruding drones. IEEE Trans. Intell. Transp. Syst. 25, 8 (Aug. 2024), 9828–9839.
@@
- Zhou et al. (2024c) Yi Zhou, Zheng Ma, Gang Liu, Zhengquan Zhang, Phee Lep Yeoh, Branka Vucetic, and Yonghui Li. 2024c. Secure multi-layer MEC systems with UAV-enabled reconfigurable intelligent surface against full-duplex eavesdropper. IEEE Trans. Commun. 72, 3 (March 2024), 1565–1577.

### v2

@@ Acknowledgements.
- Wang et al. (2019c) Peng Wang, Jiaxin Zhang, Xing Zhang, Zhi Yan, Barry G Evans, and Wenbo Wang. 2019c. Convergence of satellite and terrestrial networks: A comprehensive survey. IEEE access 8 (2019), 5550–5588.
- Wang et al. (2024c) Qixu Wang, Peng Xiao, Xiang Li, Yunxiang Qiu, Tao Zheng, and Zhiguang Qin. 2024c. Taas: Trust assessment as a service for secure communication of green edge-assisted UAV network. Peer Peer Netw. Appl. 17, 4 (July 2024), 2221–2236.
@@
- Wang et al. (2021e) Rui Wang, Xiangping Bryce Zhai, Yunlong Zhao, and Xuedong Zhao. 2021e. Delivery optimization for unmanned aerial vehicles based on minimum cost maximum flow with limited battery capacity. In Lecture Notes in Computer Science. Springer International Publishing, Cham, 77–85.
@@
- Wang et al. (2021b) Xiaoding Wang, Hui Lin, Hongyan Zhang, Dejun Miao, Qinyang Miao, and Wenxin Liu. 2021b. Intelligent drone-assisted fault diagnosis for B5G-enabled space-air-ground-space networks. IEEE Trans. Netw. Sci. Eng. 8, 4 (Oct. 2021), 2849–2860.
@@
- Wang et al. (2021c) Yuntao Wang, Zhou Su, Ning Zhang, and Dongfeng Fang. 2021c. Disaster relief wireless networks: Challenges and solutions. IEEE Wirel. Commun. 28, 5 (Oct. 2021), 148–155.
@@
- Wang et al. (2024d) Yiqian Wang, Jie Zhu, Haiping Huang, and Fu Xiao. 2024d. Bi-objective ant colony optimization for trajectory planning and task offloading in UAV-assisted MEC systems. IEEE Trans. Mob. Comput. 23, 12 (Dec. 2024), 12360–12377.
@@
- Wang et al. (2019b) Zhifang Wang, Jianguo Yu, Shangjing Lin, Junguo Dong, and Zhen Yu. 2019b. Distributed robust adaptive fault-tolerant mechanism for quadrotor UAV real-time wireless network systems with random delay and packet loss. IEEE Access 7 (2019), 134055–134062.
- Wang et al. (2021d) Zhifang Wang, Jianguo Yu, Shangjing Lin, Junguo Dong, and Zheng Yu. 2021d. Distributed robust H adaptive fault-tolerant control of amorphous and flat air-to-ground wireless self-organizing network system. Assem. Autom. 41, 6 (Nov. 2021), 641–658.
@@
- Zhang et al. (2018) Long Zhang, Zhen Zhao, Qiwu Wu, Hui Zhao, Haitao Xu, and Xiaobo Wu. 2018. Energy-aware dynamic resource allocation in UAV assisted mobile edge computing over social internet of vehicles. IEEE Access 6 (2018), 56700–56715.
- Zhang et al. (2023d) Peng Zhang, Hui Tian, Hao Luo, Xingwang Li, and Gaofeng Nie. 2023d. A hybrid fast inference approach with distributed neural networks for edge computing enabled UAV swarm. Phys. Commun. 60, 102129 (Oct. 2023), 102129.
@@
- Zhang et al. (2023c) Xuhui Zhang, Wenchao Liu, Huijun Xing, Zhenzhen Jin, Weilin Zang, Shuqiang Wang, Yanyan Shen, and Liang Xue. 2023c. Learning to hybrid offload in space-air-ground integrated mobile edge computing for IoT networks. In 2023 IEEE 13th International Conference on CYBER Technology in Automation, Control, and Intelligent Systems (CYBER) (Qinhuangdao, China). IEEE, 836–841.
@@
- Zhang et al. (2023e) Ying Zhang, Supu Xiu, Yiqing Cai, and Pengshan Ren. 2023e. Scheduling of graph neural network and Markov based UAV mobile edge computing networks. Phys. Commun. 60, 102160 (Oct. 2023), 102160.
@@
- Zheng et al. (2024) Chunxiang Zheng, Kai Pan, Jiadong Dong, Lin Chen, Qinghu Guo, Shunfeng Wu, Hanyun Luo, and Xiaolin Zhang. 2024. Multi-agent collaborative optimization of UAV trajectory and latency-aware DAG task offloading in UAV-assisted MEC. IEEE Access 12 (2024), 42521–42534.
@@
- Zhou et al. (2024c) Xunkuai Zhou, Guidong Yang, Yizhou Chen, Li Li, and Ben M Chen. 2024c. VDTNet: A high-performance visual network for detecting and tracking of intruding drones. IEEE Trans. Intell. Transp. Syst. 25, 8 (Aug. 2024), 9828–9839.